40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon May 17 19:42:11 2010, Jenne, Update, IOO, First steps toward MC mode measuring 
    Reply  Tue May 18 01:40:56 2010, Koji, Update, IOO, First steps toward MC mode measuring 
       Reply  Tue May 18 16:44:35 2010, Koji, Update, IOO, First steps toward MC mode measuring 
    Reply  Tue May 18 12:04:13 2010, rob, Update, IOO, First steps toward MC mode measuring 
Message ID: 2941     Entry time: Mon May 17 19:42:11 2010     Reply to this: 2942   2945
Author: Jenne 
Type: Update 
Category: IOO 
Subject: First steps toward MC mode measuring 

[Jenne, Kevin, Steve]

We made some progress toward getting the MC's beam profile measured.  In the end, no changes were made to anything today, but we're more prepared to go for tomorrow.

What we did:

* Grabbed the scanning slit beam scan from the PSL lab.  It's the same kind as we had here at the 40m, so Kevin was able to hook it up to the computer, and confirmed that it works.

* Opened the IOO and OMC chamber doors, and locked the MC.  Unfortunately the MC mode was awful in Yaw.  Awful like TEM(0,10+). But it still locked.  

* Confirmed that the beam went through the Faraday.  I looked at the beam before and after the Faraday on a card, and it was the same nasty beam both before and after.  So it looks like Zach did a good job aligning the Faraday and everything else.  I was going to clamp the Faraday, but I didn't yet, since I wanted to see the nice happy TEM00 mode go through without clipping before risking moving the Faraday during clamping (I don't know how heavy it is, so I'm not sure how much it might potentially move during clamping.)

* Noticed that there is a whole lot of crap on both the OMC and BS tables that's going to have to move.  In particular, one of the weights leveling the OMC table is right where I need to put MMT2.  Steve suggested putting the optic there, in its approximate place, before doing too much other stuff, since it could potentially affect the leveling of the table, and thus the input pointing to the MC.  Unfortunately, to do that I'll need to move the weight, which is definitely going to change things.  Sad face.  Moving the weight will likely be one of the first things I do tomorrow, so that all 3 profile measurements have the same configuration. 

* Before closing up, I tried to align the MC, to get back to TEM00, to no avail.  I got as far as achieving TEM11 flashing, along with a bunch of other crappy modes, but didn't get 00.  That's also on the to-do list.

What we're going to do:

* Open the chambers, and align the MC to TEM00 (using the sliders on the MC align screen).

* Check with an IR card that the beam goes through the Faraday.

* Clamp the Faraday, reconfirm.

* Remove the weight on the OMC table.

* Place MMT2 on the OMC table in it's approximate final location.

* Realign the MC, and make sure the beam goes through the Faraday.  If this doesn't happen smoothly, I may need more instruction since I've never dealt with aligning the Faraday before.  What are the appropriate mirrors to adjust? 

* Move the PZT flat steering mirror from the BS table to the IOO table.  (Thoughts on this?  This will change the table leveling, and also includes the trickiness of needing to move the connectors for the PZT.)

* Place a flat mirror on the BS table to route the MC beam out to the BS/PRM/SRM oplev table. 

* Measure the mode using the beam scan: on the BS oplev table, on the POX table, and then perhaps by shooting the beam through the beamtube on the ETMY (new convention) table.

* Place MMT1 on the BS table, use flat mirrors to get it out of the chambers, repeat measurements.

* Place MMT2 in the correct position, use flat mirrors to get it out of the chambers, repeat measurements.

All of this may require some serious cleaning-up of the BS table, which is going to be ugly, but it has to happen sometime. Hopefully I can get away with only moving a minimal number of things, in order to get these measurements done.

 

Another note: Don't trust the PSL shutter and the switch on the MEDM screens! Always use a manual block in addition!!! We discovered upon closeup that hitting the "Closed" button, while it reads back as if the shutter is closed (with the red box around the buttons), does not in fact close the shutter.  The shutter is still wide open.  This must be fixed.

ELOG V3.1.3-